# Gotta Detect ’Em All: Fake Base Station and Multi-Step Attack Detection in Cellular Networks

### v2

@@ 3.4 Proposed Solution
1 | Authentication relay attack [8] | Activity monitoring; DoS | Complete or selective DoS; Location history poisoning; Network profiling |
2 | Bidding down with AttachReject [11] | DoS | Selective DoS |
3 | Paging channel hijacking attack [8] | DoS | Complete DoS |
4 | Location tracking via measurement reports [11] | Location tracking | Leak fine-grained location |
5 | Capability Hijacking [12] | DoS, Downgrading | Selective DoS and downgrading |
6 | Incarceration with rrcReestablishReject [9] | DoS | Complete DoS |
7 | Lullaby attack using rrcReestablishRequest [9] | Battery drain | Force state change, battery draining |
8 | Bidding down with ServiceReject [11] | DoS | Selective DoS |
9 | Mobile Network Mapping (MNmap) [12] | Device Identification | Identify devices on a mobile network |
10 | Energy Depletion attack [8] | Battery drain | Battery draining |
11 | Lullaby attack with rrcResume [9] | Battery drain | Force state change, battery draining |
12 | Stealthy Kickoff Attack [8] | DoS | Complete DoS |
13 | Incarceration with rrcReject and rrcRelease [9] | DoS | Complete DoS |
14 | IMSI catching [25] | Information Leak | Leaking sensitive information |
15 | NAS counter Desynch attack [9] | DoS | Complete, prolonged DoS |
16 | X2 signalling flood [13] | Resource waste | Waste resources for the network |
17 | Handover hijacking [13] | DoS, Energy Depletion | Complete DoS and battery draining |
18 | RRC replay attack [6] | DoS | Complete DoS |
19 | Lullaby attack with rrcReconfiguration [9] | Battery drain | Force state change, battery draining |
20 | Bidding down with TAUReject [11] | DoS | Selective DoS |
21 | Panic Attack [8] | Misinformation | Artificial emergency |
@@ Appendix E Comparison with existing solutions
1 | Authentication relay attack | 46.08 | 48.49 | 2.29 | 3.14 |
2 | Bidding down with AttachReject | 52.2 | 42.56 | 2.05 | 3.19 |
3 | Paging channel hijacking attack | 51.18 | 38.6 | 3.41 | 6.81 |
4 | Location tracking via measurement reports | 51.93 | 41.52 | 2.29 | 4.26 |
5 | Capability Hijacking | 50.62 | 40.58 | 3.46 | 5.34 |
6 | Incarceration with rrcReestablishReject | 49.65 | 42.81 | 3.73 | 3.81 |
7 | Lullaby attack using rrcReestablishRequest | 44.16 | 46.36 | 3.36 | 6.12 |
8 | Bidding down with ServiceReject | 48.1 | 44.57 | 3.49 | 3.84 |
9 | Mobile Network Mapping (MNmap) | 53.05 | 40.99 | 2.09 | 3.87 |
10 | Energy Depletion attack | 52.44 | 40.95 | 2.12 | 4.49 |
11 | Lullaby attack with rrcResume | 44.79 | 46.85 | 3.89 | 4.47 |
12 | Stealthy Kickoff Attack | 52.71 | 37.84 | 4.71 | 4.74 |
13 | Incarceration with rrcReject and rrcRelease | 48.01 | 42.84 | 4.77 | 4.38 |
14 | IMSI catching | 39.17 | 53.36 | 2.69 | 4.78 |
15 | NAS counter Desynch attack | 49.81 | 42.49 | 3.04 | 4.66 |
16 | X2 signalling flood | 42.75 | 49.46 | 4.72 | 3.07 |
17 | Handover hijacking | 40.2 | 51.8 | 3.1 | 4.9 |
18 | RRC replay attack | 44.16 | 46.89 | 4.93 | 4.02 |
19 | Lullaby attack with rrcReconfiguration | 47.24 | 45.9 | 2.76 | 4.1 |
20 | Bidding down with TAUReject | 50.52 | 42.27 | 3.43 | 3.78 |
21 | Panic Attack | 50.52 | 42.27 | 3.43 | 3.78 |
@@ Appendix H List of Features
nas eps emm emc bs | lte rrc nomPDSCH RS EPRE | lte rrc dlInformationTransfer element |

### v3

@@ 3.4 Proposed Solution
| 1 | Authentication relay attack [8] | Activity monitoring; DoS | Complete or selective DoS; Location history poisoning; Network profiling |
| 2 | Bidding down with AttachReject [11] | DoS | Selective DoS |
| 3 | Paging channel hijacking attack [8] | DoS | Complete DoS |
| 4 | Location tracking via measurement reports [11] | Location tracking | Leak fine-grained location |
| 5 | Capability Hijacking [12] | DoS, Downgrading | Selective DoS and downgrading |
| 6 | Incarceration with rrcReestablishReject [9] | DoS | Complete DoS |
| 7 | Lullaby attack using rrcReestablishRequest [9] | Battery drain | Force state change, battery draining |
| 8 | Bidding down with ServiceReject [11] | DoS | Selective DoS |
| 9 | Mobile Network Mapping (MNmap) [12] | Device Identification | Identify devices on a mobile network |
| 10 | Energy Depletion attack [8] | Battery drain | Battery draining |
| 11 | Lullaby attack with rrcResume [9] | Battery drain | Force state change, battery draining |
| 12 | Stealthy Kickoff Attack [8] | DoS | Complete DoS |
| 13 | Incarceration with rrcReject and rrcRelease [9] | DoS | Complete DoS |
| 14 | IMSI catching [25] | Information Leak | Leaking sensitive information |
| 15 | NAS counter Desynch attack [9] | DoS | Complete, prolonged DoS |
| 16 | X2 signalling flood [13] | Resource waste | Waste resources for the network |
| 17 | Handover hijacking [13] | DoS, Energy Depletion | Complete DoS and battery draining |
| 18 | RRC replay attack [6] | DoS | Complete DoS |
| 19 | Lullaby attack with rrcReconfiguration [9] | Battery drain | Force state change, battery draining |
| 20 | Bidding down with TAUReject [11] | DoS | Selective DoS |
| 21 | Panic Attack [8] | Misinformation | Artificial emergency |
@@ Appendix E Comparison with existing solutions
| 1 | Authentication relay attack | 46.08 | 48.49 | 2.29 | 3.14 |
| 2 | Bidding down with AttachReject | 52.2 | 42.56 | 2.05 | 3.19 |
| 3 | Paging channel hijacking attack | 51.18 | 38.6 | 3.41 | 6.81 |
| 4 | Location tracking via measurement reports | 51.93 | 41.52 | 2.29 | 4.26 |
| 5 | Capability Hijacking | 50.62 | 40.58 | 3.46 | 5.34 |
| 6 | Incarceration with rrcReestablishReject | 49.65 | 42.81 | 3.73 | 3.81 |
| 7 | Lullaby attack using rrcReestablishRequest | 44.16 | 46.36 | 3.36 | 6.12 |
| 8 | Bidding down with ServiceReject | 48.1 | 44.57 | 3.49 | 3.84 |
| 9 | Mobile Network Mapping (MNmap) | 53.05 | 40.99 | 2.09 | 3.87 |
| 10 | Energy Depletion attack | 52.44 | 40.95 | 2.12 | 4.49 |
| 11 | Lullaby attack with rrcResume | 44.79 | 46.85 | 3.89 | 4.47 |
| 12 | Stealthy Kickoff Attack | 52.71 | 37.84 | 4.71 | 4.74 |
| 13 | Incarceration with rrcReject and rrcRelease | 48.01 | 42.84 | 4.77 | 4.38 |
| 14 | IMSI catching | 39.17 | 53.36 | 2.69 | 4.78 |
| 15 | NAS counter Desynch attack | 49.81 | 42.49 | 3.04 | 4.66 |
| 16 | X2 signalling flood | 42.75 | 49.46 | 4.72 | 3.07 |
| 17 | Handover hijacking | 40.2 | 51.8 | 3.1 | 4.9 |
| 18 | RRC replay attack | 44.16 | 46.89 | 4.93 | 4.02 |
| 19 | Lullaby attack with rrcReconfiguration | 47.24 | 45.9 | 2.76 | 4.1 |
| 20 | Bidding down with TAUReject | 50.52 | 42.27 | 3.43 | 3.78 |
| 21 | Panic Attack | 50.52 | 42.27 | 3.43 | 3.78 |
@@ Appendix H List of Features
| nas eps emm emc bs | lte rrc nomPDSCH RS EPRE | lte rrc dlInformationTransfer element | |